# Refined canonical stable Grothendieck polynomials and their duals, Part 2

### v1

@@ 1. Introduction
In the first part [10] we proved that $G_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta})$ and $g_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta})$ have Jacobi-Trudi-like formulas.
@@
Our strategy to prove Theorem 1.3 is as follows. Instead of proving (1.5) and (1.6) directly, we consider their flagged versions $G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ and $g_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ for which we can use inductive arguments in the same vein as in Wachs [19]. Using computer experiments, we were able to find suitable Jacobi–Trudi-like formulas for $G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ and $g_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ that generalize (1.3) and (1.4), respectively; see Theorem 1.4 below. Thanks to the flag conditions, $G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ and $g_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ have natural recurrences. We then prove the Jacobi–Trudi-like formulas by showing that the determinants also have the same recurrences. Although our proofs require many technical lemmas, they are elementary and we need nothing more complex than basic techniques in linear algebra such as row operations.
@@ 3. Flagged Grothendieck polynomials
Let $\lambda,\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{P}^{n}$ with $\mu\subseteq\lambda$. Suppose $r_{i}-\mu_{i}\leq r_{i+1}-\mu_{i+1},s_{i}-\lambda_{i}\leq s_{i+1}-\lambda_{i+1}+1$ and $\mu_{i}<\lambda_{i+1}$ for any $11\leq i\leq n-$ 1. If $s_{k}-\lambda_{k}+1<r_{k}-\mu_{k}$ and $\mu_{k}<\lambda_{k}$ for some $1\leq k\leq n$, then
@@
$\tilde{G}_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=\tilde{G}_{({\lambda}^{(1)})^{\prime}/({\mu}^{(1)})^{\prime}}^{\operatorname{col}({\bm{r}}^{(1)},{\bm{s}}^{(1)})}(\bm{x};\bm{\alpha},\bm{\beta})\\ \times\phi_{\bm{\alpha}}^{k-1}\left(\tilde{G}_{((\lambda_{k}-1))^{\prime}/((\mu_{k}))^{\prime}}^{\operatorname{col}((r_{k}),(s_{k}-1))}(\bm{x};\bm{\alpha},\bm{\beta})\right)\phi_{\bm{\alpha}}^{k}\left(\tilde{G}_{({\lambda}^{(2)})^{\prime}/({\mu}^{(2)})^{\prime}}^{\operatorname{col}({\bm{r}}^{(2)},{\bm{s}}^{(2)})}(\bm{x};\bm{\alpha},\bm{\beta})\right),$
@@
(3.19) $\tilde{G}_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=\tilde{G}_{({\lambda}^{(1)})^{\prime}/({\mu}^{(1)})^{\prime}}^{\operatorname{col}({\bm{r}}^{(1)},{\bm{s}}^{(1)})}(\bm{x};\bm{\alpha},\bm{\beta})\phi_{\bm{\alpha}}^{k}\left(\tilde{G}_{({\lambda}^{(2)})^{\prime}/({\mu}^{(2)})^{\prime}}^{\operatorname{col}({\bm{r}}^{(2)},{\bm{s}}^{(2)})}(\bm{x};\bm{\alpha},\bm{\beta})\right).$
The same argument also shows that $\tilde{G}_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ is equal to the right-hand side of (3.19). Therefore we obtain (3.17). ∎
@@ 4. Flagged dual Grothendieck polynomials
$\operatorname{wt}(T(i,j))=\begin{cases}-\alpha_{j}&\mbox{if $T(i,j)$ is marked,}\\ \beta_{i-1}&\mbox{if $T(i,j)$ is not marked and $T(i,j)=T(i-1,j)$,}\\ x_{T(i,j)}&\mbox{if $T(i,j)$ is not marked and $T(i,j)\neq T(i-1,j)$.}\end{cases}$
@@
We now suppose $i<j$. If $\mu_{j}<\lambda_{i}$, then by the same argument in (4.4), we have $R_{i,j}=0$. This is a contradiction to the assumption $L_{i,j}\neq R_{i,j}$, so we must have $\lambda_{i}\leq\mu_{j}$. Then, since $\mu_{i}\leq\lambda_{i}\leq\mu_{j}\leq\mu_{i}$, we also have $\mu_{i}=\lambda_{i}$. For integers $p$ and $q$ satisfying $1\leq p\leq i\leq q\leq n$, we have $\lambda_{q}-\mu_{p}+p-q\leq 0$ and the equality holds if and only if $p=q=i$, which means that $L_{q,p}=R_{q,p}=\chi(p=q=i)$. From this, the matrix $R$ is a block upper triangular matrix and it gives $\det R=\det(R_{k,l})_{k,l=1}^{i-1}\det(R_{k,l})_{k,l=i+1}^{n}$. This means that the $(i,j)$-entry of $R$ does not contribute to the determinant of the matrix $R$, which completes the proof. ∎
@@
Let $\lambda$ be a dented partition and $\mu$ a partition with $\mu\subseteq\lambda$. Let $\bm{r},\bm{s}\in\mathbb{P}^{n}$ and $I\subseteq[n]$. If $\bm{r}\not\leq\bm{s}$, we define ${\operatorname{MRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}(\lambda/\mu)=\emptyset$. If $\bm{r}\leq\bm{s}$, we define ${\operatorname{MRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}({\lambda/\mu})$ to be the set of RPPs $T$ of shape ${\lambda/\mu}$ in which every entry $T(i,j)$ with $T(i,j)=T(i,j+1)$ may be marked, where for each $i\in[n]$ we define
@@
$\operatorname{wt}(T(i,j))=\begin{cases}-\alpha_{j}&\mbox{if $T(i,j)$ is marked,}\\ \beta_{i-1}&\mbox{if $T(i,j)$ is not marked and $T(i,j)=T(i-1,j)$,}\\ x_{T(i,j)}&\mbox{if $T(i,j)$ is not marked and $T(i,j)\neq T(i-1,j)$.}\end{cases}$
@@
If $F=M$, then the two statements follow immediately from the definition of $M(\lambda,\mu,\bm{r},\bm{s},I)$. For the case $F=D$, let
@@
We proceed by induction on $(M,N)$, where $M:=\sum_{i=1}^{n}(s_{i}-r_{i})$ and $N:=|{\lambda/\mu}|$. There are two base cases: the case $M<0$ and the case $N=0$, which can be easily obtained from Lemmas 4.15 and 4.16.
@@
$\operatorname{wt}_{R}(T(i,j))=\begin{cases}-\alpha_{j-1}&\mbox{if$T(i,j)$ is marked,}\\ \beta_{i}&\mbox{if $T(i,j)$ is not marked and $T(i,j)=T(i+1,j)$,}\\ x_{T(i,j)}&\mbox{if $T(i,j)$ is not marked and $T(i,j)\neq T(i+1,j)$.}\end{cases}$
@@
$\operatorname{wt}_{B}(T(i,j))=\begin{cases}-\alpha_{i-1}&\mbox{if$T(i,j)$ is marked,}\\ \beta_{j}&\mbox{if $T(i,j)$ is not marked and $T(i,j)=T(i,j+1)$,}\\ x_{T(i,j)}&\mbox{if $T(i,j)$ is not marked and $T(i,j)\neq T(i,j+1)$.}\end{cases}$
@@
Let $\lambda$ be a dented partition and $\mu$ a partition with $\mu\subseteq\lambda$. Let $\bm{r},\bm{s}\in\mathbb{P}^{n}$ and $I\subseteq[n]$. If $\bm{r}\not\leq\bm{s}$, we define ${\operatorname{BMRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}(\lambda/\mu)=\emptyset$. If $\bm{r}\leq\bm{s}$, we define ${\operatorname{BMRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}({\lambda/\mu})$ to be the set of RPPs $T$ of shape ${\lambda/\mu}$ in which every entry $T(i,j)$ with $T(i,j)=T(i-1,j)$ may be marked. For each $i\in[n]$, we define
@@
$\operatorname{wt}_{B}(T(i,j))=\begin{cases}-\alpha_{i-1}&\mbox{if$T(i,j)$ is marked,}\\ \beta_{j}&\mbox{if $T(i,j)$ is not marked and $T(i,j)=T(i,j+1)$,}\\ x_{T(i,j)}&\mbox{if $T(i,j)$ is not marked and $T(i,j)\neq T(i,j+1)$.}\end{cases}$
@@
The case $F=M$ is immediate from the definition of $M^{\prime}(\lambda,\mu,\bm{r},\bm{s},I)$, so we only consider the case $F=D$.
@@ Acknowledgments
The authors are grateful to Travis Scrimshaw for many helpful comments and fruitful discussions. They also thank Darij Grinberg for useful comments.

### v2

@@ 1. Introduction
In the first part [10] we proved that $G_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta})$ and $g_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta})$ have Jacobi–Trudi-like formulas.
@@
Our strategy to prove Theorem 1.3 is as follows. Instead of proving (1.5) and (1.6) directly, we consider their flagged versions $G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ and $g_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$, see Definitions 3.5 and 4.4, for which we can use inductive arguments in the same vein as in Wachs [19].
Using computer experiments, we were able to find suitable Jacobi–Trudi-like formulas for $G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ and $g_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ that generalize (1.3) and (1.4), respectively; see Theorem 1.4 below. Thanks to the flag conditions, $G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ and $g_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ have natural recurrences. We then prove the Jacobi–Trudi-like formulas by showing that the determinants also have the same recurrences. Although our proofs require many technical lemmas, they are elementary and we need nothing more complex than basic techniques in linear algebra such as row operations.
@@ 3. Flagged Grothendieck polynomials
where $\gamma^{(1)}=(\gamma_{1},\dots,\gamma_{k})$ and $\gamma^{(2)}=(\gamma_{k+1},\dots,\gamma_{n})$ for each $\gamma\in\{\lambda,\mu,\bm{r},\bm{s}\}$.
@@
Let $\lambda,\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{P}^{n}$ with $\mu\subseteq\lambda$. Suppose $r_{i}-\mu_{i}\leq r_{i+1}-\mu_{i+1},s_{i}-\lambda_{i}\leq s_{i+1}-\lambda_{i+1}+1$. If $s_{k}-\lambda_{k}+1<r_{k}-\mu_{k}$ and $\mu_{k}<\lambda_{k}$ for some $1\leq k\leq n$, then
@@
$\tilde{G}_{(\lambda-\epsilon_{k})^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=\tilde{G}_{({\lambda}^{(1)})^{\prime}/({\mu}^{(1)})^{\prime}}^{\operatorname{col}({\bm{r}}^{(1)},{\bm{s}}^{(1)})}(\bm{x};\bm{\alpha},\bm{\beta})\\ \times\phi_{\bm{\alpha}}^{k-1}\left(\tilde{G}_{((\lambda_{k}-1))^{\prime}/((\mu_{k}))^{\prime}}^{\operatorname{col}((r_{k}),(s_{k}-1))}(\bm{x};\bm{\alpha},\bm{\beta})\right)\phi_{\bm{\alpha}}^{k}\left(\tilde{G}_{({\lambda}^{(2)})^{\prime}/({\mu}^{(2)})^{\prime}}^{\operatorname{col}({\bm{r}}^{(2)},{\bm{s}}^{(2)})}(\bm{x};\bm{\alpha},\bm{\beta})\right),$
@@
(3.19) $\tilde{G}_{(\lambda-\epsilon_{k})^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=\tilde{G}_{({\lambda}^{(1)})^{\prime}/({\mu}^{(1)})^{\prime}}^{\operatorname{col}({\bm{r}}^{(1)},{\bm{s}}^{(1)})}(\bm{x};\bm{\alpha},\bm{\beta})\phi_{\bm{\alpha}}^{k}\left(\tilde{G}_{({\lambda}^{(2)})^{\prime}/({\mu}^{(2)})^{\prime}}^{\operatorname{col}({\bm{r}}^{(2)},{\bm{s}}^{(2)})}(\bm{x};\bm{\alpha},\bm{\beta})\right).$
The same argument also shows that $\tilde{G}_{(\lambda-\epsilon_{k})^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})$ is equal to the right-hand side of (3.19). Therefore we obtain (3.17). ∎
@@ 4. Flagged dual Grothendieck polynomials
$\operatorname{wt}(T(i,j))=\begin{cases}-\alpha_{j}&\mbox{if $T(i,j)$ is marked,}\\ \beta_{i-1}&\mbox{if $T(i,j)$ is not marked and $T(i,j)=T(i-1,j)$,}\\ x_{T(i,j)}&\mbox{otherwise.}\end{cases}$
@@
We now suppose $i<j$. If $\mu_{j}<\lambda_{i}$, then by the same argument in (4.4), we have $R_{i,j}=0$. This is a contradiction to the assumption $L_{i,j}\neq R_{i,j}$, so we must have $\lambda_{i}\leq\mu_{j}$. Then, since $\mu_{i}\leq\lambda_{i}\leq\mu_{j}\leq\mu_{i}$, we also have $\mu_{i}=\lambda_{i}$. Thus by 2.5, we obtain $\det R=\det(R_{k,l})_{k,l=1}^{i-1}\det(R_{k,l})_{k,l=i+1}^{n}$. This means that the $(i,j)$-entry of $R$ does not contribute to the determinant of the matrix $R$, which completes the proof. ∎
@@
Let $\lambda\in\operatorname{DPar}_{n}$ and $\mu\in\operatorname{Par}_{n}$ with $\mu\subseteq\lambda$, and let $\bm{r},\bm{s}\in\mathbb{P}^{n}$ and $I\subseteq[n]$. If $\bm{r}\not\leq\bm{s}$, we define ${\operatorname{MRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}(\lambda/\mu)=\emptyset$. If $\bm{r}\leq\bm{s}$, we define ${\operatorname{MRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}({\lambda/\mu})$ to be the set of RPPs $T$ of shape ${\lambda/\mu}$ satisfying the following conditions.
  • For all $(i,j)\in{\lambda/\mu}$, we have $r_{i}\leq T(i,j)\leq s_{i}$.
  • If $(k,\lambda_{k})$ is the minimal cell of $\lambda$ and $k-1\in I$, then $T(k,\lambda_{k})\geq s_{k-1}$.
  • Every entry $T(i,j)$ with $T(i,j)=T(i,j+1)$ may be marked, where for each $i\in[n]$ we define
@@
$\operatorname{wt}(T(i,j))=\begin{cases}-\alpha_{j}&\mbox{if $T(i,j)$ is marked,}\\ \beta_{i-1}&\mbox{if $T(i,j)$ is not marked and $T(i,j)=T(i-1,j)$,}\\ x_{T(i,j)}&\mbox{otherwise.}\end{cases}$
Here, the equality $T(i,j)=T(i-1,j)$ means that their underlying integers are equal.
@@
We first consider the case $F=M$. For the first statement, suppose that $\lambda$ is not a partition and $s_{k-1}=s_{k}$. Let $T\in{\operatorname{MRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}({\lambda/\mu})$. Note that since $k-1\in I$, we have $T(k-1,\lambda_{k-1}+1)=s_{k-1}$ and $s_{k-1}\leq T(k,\lambda_{k})\leq s_{k}$. Thus $T(k-1,\lambda_{k})=T(k-1,\lambda_{k-1}+1)=s_{k-1}=s_{k}$ and $T(k,\lambda_{k})=s_{k}$. This implies that $\operatorname{wt}(T(k,\lambda_{k}))$ is $-\alpha_{\lambda_{k}}$ or $\beta_{k-1}$ if $k\in I$ and $\beta_{k-1}$ otherwise. Let $T^{\prime}$ be the tableau obtained from $T$ by removing $T(k,\lambda_{k})$. Then $T\mapsto T^{\prime}$ is a bijection from $M(\lambda,\mu,\bm{r},\bm{s},I)$ to $M(\lambda-\epsilon_{k},\mu,\bm{r},\bm{s},I\cup\{k\})$ such that $\operatorname{wt}(T)=\operatorname{wt}(T^{\prime})\cdot(\beta_{k-1}-\chi(k\in I)\alpha_{\lambda_{k}})$. Therefore,
$M(\lambda,\mu,\bm{r},\bm{s},I)=(\beta_{k-1}-\chi(k\in I)\alpha_{\lambda_{k}})M(\lambda-\epsilon_{k},\mu,\bm{r},\bm{s},I\cup\{k\}),$
which is the first statement. The second statement can be proved similarly with two cases $T(k,\lambda_{k})=s_{k}$ and $T(k,\lambda_{k})<s_{k}$, which give $(x_{s_{k}}-\chi(k\in I)\alpha_{\lambda_{k}})F(\lambda-\epsilon_{k},\mu,\bm{r},\bm{s},I\cup\{k\})$ and $F(\lambda,\mu,\bm{r},\bm{s}-\epsilon_{k},I\setminus\{k\})$, respectively.
Now we consider the case $F=D$. Let
@@
We proceed by induction on $(M,N)$, where $M:=\sum_{i=1}^{n}(s_{i}-r_{i})$ and $N:=|{\lambda/\mu}|$. There are two base cases: the case $M<0$, which follows from Lemmas 4.15 and 4.17, and the case $N=0$, which follows from Lemma 4.16.
@@
$\operatorname{wt}_{R}(T(i,j))=\begin{cases}-\alpha_{j-1}&\mbox{if$T(i,j)$ is marked,}\\ \beta_{i}&\mbox{if $T(i,j)$ is not marked and $T(i,j)=T(i+1,j)$,}\\ x_{T(i,j)}&\mbox{otherwise.}\end{cases}$
@@
$\operatorname{wt}_{B}(T(i,j))=\begin{cases}-\alpha_{i-1}&\mbox{if$T(i,j)$ is marked,}\\ \beta_{j}&\mbox{if $T(i,j)$ is not marked and $T(i,j)=T(i,j+1)$,}\\ x_{T(i,j)}&\mbox{otherwise.}\end{cases}$
@@
Let $\lambda\in\operatorname{DPar}_{n}$ and $\mu\in\operatorname{Par}_{n}$ with $\mu\subseteq\lambda$, and let $\bm{r},\bm{s}\in\mathbb{P}^{n}$ and $I\subseteq[n]$. If $\bm{r}\not\leq\bm{s}$, we define ${\operatorname{BMRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}(\lambda/\mu)=\emptyset$. If $\bm{r}\leq\bm{s}$, we define ${\operatorname{BMRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}({\lambda/\mu})$ to be the set of RPPs $T$ of shape ${\lambda/\mu}$ satisfying the following conditions.
  • For all $(i,j)\in{\lambda/\mu}$, we have $r_{i}\leq T(i,j)\leq s_{i}$.
  • If $(k,\lambda_{k})$ is the minimal cell of $\lambda$ and $k-1\in I$, then $T(k,\lambda_{k})\geq s_{k-1}$.
  • Every entry $T(i,j)$ with $T(i,j)=T(i-1,j)$ may be marked, where for each $i\in[n]$ we define
@@
$\operatorname{wt}_{B}(T(i,j))=\begin{cases}-\alpha_{i-1}&\mbox{if$T(i,j)$ is marked,}\\ \beta_{j}&\mbox{if $T(i,j)$ is not marked and $T(i,j)=T(i,j+1)$,}\\ x_{T(i,j)}&\mbox{otherwise.}\end{cases}$
@@
The case $F=M$ can be proved similarly as in the proof of Lemma 4.18. Hence, we only consider the case $F=D$.
@@ Acknowledgments
The authors thank Travis Scrimshaw for many helpful comments and fruitful discussions, and Darij Grinberg for valuable feedback. They are also grateful to the anonymous referees for their careful reading and thoughtful suggestions.
The authors were supported by the National Research Foundation of Korea (NRF) grants #2022R1A2C101100911 and #2016R1A5A10080. Byung-Hak Hwang was supported by a KIAS Individual Grant (MG098201) at Korea Institute for Advanced Study. Minho Song was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2022R1C1C2009025).
Appendix A Marked rim border tableaux
In this appendix, we introduce marked rim border tableaux and provide a weight-preserving bijection between marked rim border tableaux and marked RPPs. This shows that Theorem 4.2 generalizes Yeliussizov’s combinatorial model for the dual Grothendieck polynomial $g^{(\alpha,\beta)}_{\lambda}(\bm{x})$.
Definition A.1.
For a reverse plane partition $T$, let $T_{i}$ denote the part of $T$ consisting of cells with entry $i$. Define the border $R_{i}$ of $T_{i}$ as the part consisting of cells $c$ where the upper-left cell of $c$ is not in $T_{i}$. The inner part of $T_{i}$ is defined to be $T_{i}\setminus R_{i}$. Each border $R_{i}$ can be partitioned by cutting along a vertical side of a cell. The resulting cut blocks are called rim hooks. A marked rim border tableau is a reverse plane partition $T$ in which for each $i$, the border $R_{i}$ of $T_{i}$ is partitioned into rim hooks and each cell in the inner part $T_{i}\setminus R_{i}$ can be marked. We denote by $\operatorname{MRBT}({\lambda/\mu})$ the set of marked rim border tableaux of shape ${\lambda/\mu}$.
Definition A.2.
For a marked rim border tableau $T$, we define the weight $\operatorname{wt}(T)$ to be the product of $\operatorname{wt}(c)$ for all cells $c$ in $T$, where $\operatorname{wt}(c)$ is defined as follows.
  • If $c$ is in a rim hook $R$ with entry $i$, then
    $\operatorname{wt}(c)=\begin{cases}\alpha&\mbox{ if the right cell of $c$ is in$R$, }\\ \beta&\mbox{ if the upper cell of $c$ is in $R$, }\\ x_{i}&\mbox{ otherwise. }\end{cases}$
  • If $c$ is in an inner part, then
    $\operatorname{wt}(c)=\begin{cases}\alpha&\mbox{ if $c$ is marked, }\\ \beta&\mbox{ if $c$ is not marked. }\end{cases}$
Yeliussizov’s result can be restated as follows.
Theorem A.3.
[…]
$g_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})=\sum_{T\in\operatorname{MRBT}(\lambda)}\operatorname{wt}(T).$
The following proposition with $\mu=\emptyset$ shows that Theorem 4.2 generalizes the above result of Yeliussizov.
Proposition A.4.
[…]
Proof.
We establish a bijection $\varphi:\operatorname{MRBT}({\lambda/\mu})\to{\operatorname{MRPP}}({\lambda/\mu})$ as follows. Let $T\in\operatorname{MRBT}({\lambda/\mu})$. As before, let $T_{i}$ be the part of $T$ consisting of cells with entry $i$, and let $R_{i}$ be its border.
For each $i$, we shift all cells in the inner part $T_{i}\setminus R_{i}$ to the left by one. Next, we move each overlapping cell in $R_{i}$ to the rightmost cell in the same row in $T_{i}$. Then we assign a mark to each cell with $\alpha$. We define $\varphi(T)$ to be the resulting tableau; see Figure 8.
Note that all overlapping cells in $R_{i}$ in the process are in a rim border with weight $\beta$. Thus, the resulting tableau $\varphi(T)$ is a marked RPP. One can see that this gives a weight-preserving bijection. ∎
$x_{i}$$\beta$$\alpha$$x_{i}$$\alpha$$\beta$$x_{i}$$\alpha$$\beta$$\beta$$\alpha$$\alpha$$\beta$$\alpha$$\beta$$\beta$$\alpha$$\beta$$\alpha$$\alpha$$\alpha$$\beta$ $x_{i}$$\beta$$\alpha$$x_{i}$$\alpha$$\beta$$x_{i}$$\alpha$$\beta$$\beta$$\alpha$$\alpha$$\alpha$$\beta$$\beta$$\alpha$$\beta$$\alpha$$\alpha$$\alpha$$\beta$$\beta$
Figure 8. The left tableau shows the part $T_{i}$ in a rim border tableau $T$, where the border is gray and the inner part is white. Each inner cell has weight $\alpha$ if marked, $\beta$ otherwise. The right shows the corresponding part with entry $i$ in the marked RPP $\varphi(T)$, where the marked cells are those containing an $\alpha$.